# Towards Contextual Sensitive Data Detection

### v1

@@ 8 Conclusion
The rise of open data portals are key in facilitating the democratization of access to important data. This development, while positive, comes with risks of having sensitive data being used for harmful purposes by malicious actors as well as indirect leakage into language models which are fueled by any data on the web. Protecting sensitive data from getting published through such portals, by accident, is therefore key. In this paper, we first revisit the concept of sensitive data which is typically focused on personal-level data like personal-identifiable information (PII). Methods for detecting sensitive data are limited: they are too specific to cover all sensitive data, and too generic to only flag data that is actually sensitive. We have introduced a broader, contextual, perspective on sensitive data, and present a framework that formalizes contextual sensitive data from two angles. Experiments show that type-contextualization through an LLM-assisted mechanism that first detects the type of the data values (e.g. names) followed by a reflection on their sensitivity, significantly improves precision and recall over commercial tools and out-of-the-box LLM baselines. A case study with humanitarian datasets on the Humanitarian Data Exchange shows that the domain-contextualization through automated retrieval of rule-based heuristics from domain-specific documents effectively grounds an LLM’s assessment of data sensitivity. With our conceptualization of contextual sensitive data, we intend to encourage researchers and practitioners to consider a broader view on data sensitivity. We believe that our framework for detecting sensitive data provides a stepping stone in this important direction.
@@ Acknowledgements
We are grateful for the collaboration with the Responsible Data team of the United Nations OCHA Humanitarian Data Centre. Specifically, we like to thank Alexandru Gartner, Godfrey Takavarasha, Jos Berens, Javier Téran, Metasebya Sahlu, Melanie Rabier, Nafissatou Pouye, Sarah Telford, Serban Teodorescu, and Shahrooz Badkoubei, for their invaluable contributions to this project. This work has been supported by funding from an AiNed grant from NWO (NGF.1607.22.045), a grant from SAP, and credit grants from OpenAI, Google and Cohere.

### v2

@@ 7 Conclusion
With the rise of open data portals to democratize data access, comes the risk of exposing sensitive data to malicious actors or language models fueled by public data. Protecting sensitive data is therefore key, but methods for this are limited: they are too specific to capture all sensitive data, and too generic to only suppress data that is actually sensitive. We argue for contextualizing sensitive data detection along two dimensions. First, type contextualization, which considers fine-grained aspects beyond static sensitivity types. Our detect-then-reflect mechanism shows that first detecting if data values belong to a sensitive type, then reflecting on their true sensitivity, increases recall over existing tools with up to 49%. Second, we introduce domain contextualization which considers domain-specific factors external to the data in order to assess its sensitivity. A case study with humanitarian datasets highlights that contextualizing the sensitivity assessment in relevant domain-specific rules through a retrieval stage, extends the type coverage and enables domain-grounded explanations. We believe that our work provides a stepping stone in this important direction.
Limitations and Risks
Evaluation coverage
In our humanitarian case study, we show that domain contextualization leads to higher precision while preserving coverage, and also provides useful grounded explanations to data quality reviewers. The limited availability of labeled data in this case study, might affect generalization of the insights for domain-contextualized sensitive data detection. Moreover, while the conceptual framework and mechanisms are modality agnostic (Section 3), the lack of datasets with labeled sensitivity and the data setting of the humanitarian case study have resulted in evaluations for tabular data only, for which we constructed and publish hand-labeled datasets. We call for more diverse labeled datasets, to advance this important problem, and evaluate the framework and mechanisms for different modalities, languages, and domains.
Modality coverage
Despite the focus of our implementation and evaluation on tabular data. The proposed mechanisms easily extend to other modalities, such as free-form text data: 1) type-informed sensitivity in free-form text can be implemented efficiently by concatenating a decomposition component with a more lightweight learned type predictor Hulsebos et al. (2019) and performing reflection in batches, 2) for domain-contextualization of free-form text, a hierarchical reflection procedure could be efficient. For example, starting with assessing high-level segments of the text (document-level, paragraph-level, sentence-level, etc Lin et al. (2025)), and incrementally narrowing down.
Sensitivity specification dependency
The presented mechanisms for detecting sensitive data can enable automation but also assist humans in data reviewing by providing context-grounded explanations with explicit citations to relevant sensitivity rules. For successful adoption in practice, however, the mechanisms depend on having access to reasonable sensitivity instructions in the form of 1) fine-grained semantic types of data that are deemed sensitive for type contextualization, and 2) reasonably specified assessment rules in domain-specific documents or (web) data sources.
Data access delay and further automation
A potential negative consequence of automated data sensitivity detection is that critical datasets might be blocked that are time-sensitive and could be instrumental for solving societal problems, for example, in humanitarian domains. The mechanisms could be tailored around the desired precision/recall trade-off to reduce false alarms. In this work, we only present mechanisms for detecting contextual data sensitivity, and not for automated remedies. In practice, a report could be generated based on existing cases to instruct the user on how to remedy the sensitive data, or potentially even automatically remedy this, which is also an important direction for future research.
@@ Acknowledgements
We are grateful for the collaboration with the Responsible Data team of the United Nations OCHA Humanitarian Data Centre. Specifically, we thank Alexandru Gartner, Godfrey Takavarasha, Jos Berens, Javier Téran, Metasebya Sahlu, Melanie Rabier, Nafissatou Pouye, Sarah Telford, Serban Teodorescu, and Shahrooz Badkoubei, for their invaluable contributions to this project. This work has been supported by funding from an AiNed grant from NWO (NGF.1607.22.045) and credit grants from OpenAI.